A series R L circuit is connected to a 110V ac source. If the voltage across the resistor is 85 V, find the power dissipated in the circuit


1
Expert's answer
2021-05-28T02:26:51-0400

"P_{avr}=\frac{1}{2}I_0U_R=\frac{1}{2}\frac{U_R^2}{R}"


Assume that "R=20\ \Omega" . So, we have


"P_{avr}=\frac{1}{2}\frac{U_R^2}{R}=\frac{1}{2}\frac{85^2}{20}=180.6\ (W)" . Answer